OpenStudy (tlynne16):

Lane Manufacturing estimates that its weekly profit, P, in thousands of dollars, can be approximated by the formula P = -0.1x2 + 0.6x + 3 where x is the number of units produced per week, in thousands. a. How many units should the company produce per week to earn the maximum profit?

OpenStudy (tlynne16):

@Mehek14

OpenStudy (fwizbang):

You could rewrite the equation as P=a(x-h)^2+k. Can you find a,h, and k? (I'm assuming this is an algebra question, not calculus)

OpenStudy (photon336):

@Tlynne16 can you use calculus for this or not? it's pretty straightforward that way. Hey @fwizbang how would we go about finding the maximum value? \[P = a(x-h)^{2}+k \]

OpenStudy (fwizbang):

If a<0, then P has a maximum at h=k.

OpenStudy (fwizbang):

err...x=h.

OpenStudy (tlynne16):

what is the easiest way to find this?

OpenStudy (tlynne16):

it is an algebra question

OpenStudy (fwizbang):

So, if p=a(x-h)^2+k, then P= ax^2-2ahx +ah^2+k. Compare this to your equation. Whatever multiplies x^2 is a, whatever multiplies x is -2ah. So, a=? -2ah=?

OpenStudy (tlynne16):

I am so confused.

OpenStudy (photon336):

@Tlynne16 try to compare coefficients

OpenStudy (fwizbang):

The equation you are given is \[P=-0.1x^2+6x+3\] In this equation, what multiplies \[x^2\]?